OpenStudy (anonymous):

integral using partial fractions(2x+4)/x^3-2x^2 dx

OpenStudy (anonymous):

\[\int\limits_{}^{} 2x+4\div \sqrt{x ^{3}-2x ^{2}}dx\]

OpenStudy (anonymous):

the square root isnt in the problem im sorry i looked at the wrong problem

OpenStudy (anonymous):

Split into 2 parts first!

OpenStudy (anonymous):

thats what i was thinking

OpenStudy (anonymous):

Did you see x canceled out in the first part?

OpenStudy (anonymous):

nah i think splits up in to 3 parts and the last one you use long divsion to solve

OpenStudy (anonymous):

but im not sure im just trying to match it up to an example i had

OpenStudy (anonymous):

We're not in the same pace! I'm talking about split the upper! then from the second part we split the lower into 3 parts.

OpenStudy (anonymous):

The one split in 3 parts apply partial fractions technique!

OpenStudy (anonymous):

It's fine that you just go straight split the lower into 3 parts!

OpenStudy (anonymous):

2x + 4 = A/x + B/x² + C/ ( x -2)

OpenStudy (anonymous):

Denominator = x² ( x-2)

OpenStudy (anonymous):

would it come out to be \[2x+4/x -\ln \left| x \right|+2\ln \left[ x-1 \right]+C\]

OpenStudy (anonymous):

Let me check!

OpenStudy (anonymous):

= 2 ln|x| + 2/x - 2ln|x-2| + C

OpenStudy (anonymous):

So, you're okie with the result?

OpenStudy (anonymous):

i was off how did i mess up my numbers?

OpenStudy (anonymous):

Just take time to familiarize your self with the form!

OpenStudy (anonymous):

yeah ill look over it again thanks!

OpenStudy (anonymous):

practice makes perfect :)